OpenStudy (anonymous):

Simplify using your knowledge of properties. 2(a + 4) + -2a + -8

OpenStudy (akashdeepdeb):

2(a+4) + -2(a+4) [WE CAN GROUP THE SECOND PART OF THE EXPRESSION TOGETHER] Now as they are the same in magnitude but are +ve and -ve [THEY get cut] Thus the answer is ZERO (0) :)

OpenStudy (akashdeepdeb):

IN -2a-8 we took -2 commom outside so it became -2(a+4) We can take (a+4) common and thus right it as (a+4)(2-2) = (a+4)0 = 0 :)